OpenStudy (anonymous):

how to find slope and y-intercept of 2y=-3x+6

OpenStudy (anonymous):

y=mx+c m-slope c-y-intercept so slope=-3/2 y-intercept=3

OpenStudy (aroub):

y=mx+b y=-3/2x+3 slope=-3/2 y intercept =3

OpenStudy (anonymous):

thank guys

OpenStudy (asadkarim7):

to find the slope we have to make it in the form y= mx +c we would divde both sides with 2 2y =-3x +6 y = -3/2 + 6/2 therefore m = - 3/2 c = 6/2 = 3